Circular No. 8061

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003W IN UGC 5234 Further to IAUC 8058, W. Li and D. Weisz report the LOTOSS dis-

covery of a supernova on unfiltered KAIT images taken on Jan. 28.4 (mag ∼ 17.3) and 29.4 UT (mag ∼ 16.9), confirmed on Jan. 29.4 (mag ∼ 17.0) with the Tenagra II telescope by M. Schwartz. SN 2003W is located at α = 9^{h}46^{m}49^{s}.48, δ = +16^{o}02′37′′.6 (equinox 2000.0), which is 0′′.3 west and 3′′.4 north of the nucleus of UGC 5234. A KAIT image taken on Jan. 16.4 showed nothing at this position (limiting mag ∼ 19.0).

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 320–880 nm) of SN 2003W, obtained by Matheson on Jan. 29.34 UT with the MMT 6.5-m telescope (+ Blue Channel spectrograph), shows it to be a type-Ia supernova well before maximum. The spectrum consists of a blue continuum with well-defined Si II and Ca II features, similar to that of SN 1994D at 12 days before maximum (Filippenko 1997, ARAA 35, 309). Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 6017 km/s for the host galaxy, the supernova expansion velocity is 20200 km/s for Si II (rest 635.5 nm) and 27600 km/s for Ca II (rest 395.1 nm). These expansion velocities imply an age of ≈ 2 weeks before maximum, based upon comparison with values presented by Jha et al. (1999, Ap.J. Supp. 125, 73).

SUPERNOVA 2003U IN NGC 6365A Matheson et al. add that a spectrum (range 320–880 nm) of SN 2003U

(cf. IAUC 8059), obtained as above on Jan. 29.53 UT, shows it to be a type-Ia supernova. Adopting the NED recession velocity of 8496 km/s for the host galaxy, the supernova expansion velocities are 13300 km/s for Si II (rest 635.5 nm) and 19200 km/s for Ca II (rest 395.1 nm). The spectralfeature age of the supernova is 5 ± 2 days before maximum light.

SUPERNOVA 2003L IN NGC 3506 S. R. Kulkarni, California Institute of Technology, reports that A. M.

Soderberg, E. Berger, D. A. Frail, and he have observed a bright, transient radio source at the optical position of SN 2003L (IAUC 8048) with the Very Large Array (fluxes: Jan. 26.23 UT, 1.4 ± 0.03 mJy at 8.5 GHz; Jan. 28.22, 1.6 ± 0.06 mJy at 8.5 GHz and 3.2 ± 0.08 mJy at 22 GHz). They comment: “The radio luminosity is high, being 20 percent that of SN 1998bw. Alternatively, the circumstellar medium could be quite dense.”

2003 January 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green